# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v24

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.6.0$.
@@ 1 Introduction
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10, 11]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [12, 13, 14, 15, 16, 17, 18]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [19, 20]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [21]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22]. An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [23] using the SinMin Sinhala corpus created by Upeksha et al. [24, 25] is given in Appendix B.
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [28] and Wijeratne et al. [35]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [45, 46] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [47], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.6.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR. Given that this work was initially published in 2019, when citing this work, it is recommended to cite the most recent version to avoid the reviewers questioning its continued contemporary relevance. An example of an updated reference is given in Code Block LABEL:lst:ref.
@@
journal={arXiv preprint arXiv:1906.02358v24},
@@
Figure 8 in Appendix C shows the probability of studies from the institutions to which the most prolific authors from Figure 7 are affiliated citing institutions of the same set. We observe both interesting and disturbing trends which we discuss in the figure caption. Figure 9 in Appendix C shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper. If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication. For example, the paper Building a wordnet for Sinhala [48] is counted for both Nisansa de Silva and Gihan Dias. If a single paper contributes to more than one research area, that paper is counted for all of the research areas to which it contributes. For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [28] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
The remainder of this survey is organized as follows; Section 2 introduces some important properties and conventions of the Sinhala language which are important for the development and understanding of Sinhala NLP. Section 3 discusses the various tools and research available for Sinhala NLP. In this section we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work. We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka. Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work. Finally, Section 5, concludes the survey.
@@ 3.2 Data Sets
Singh et al. [159] introduced the Aya Data set [^87] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^88] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^89] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality. Kudugunta et al. [160] introduced MADLAD-400909090https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents. They have also released the trained checkpoints on GitHub [^91] https://github.com/google-research/google-research/tree/master/madlad_400. A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al. [161] along with data [^92] https://huggingface.co/datasets/afaji/cvqa and a leader board [^93] https://eval.ai/web/challenges/challenge-page/2305/overview. Pushpananda et al. [162] claims to have created a manually annotated parallel corpus named TamSiPara with 25k parallel sentences in the Si-Ta direction and 22k parallel sentences in the Ta-Si direction. But it is not freely and openly accessible. While the stated purpose of Ranasinghe [163] is creating a tokenizer for Sinhala and not creating a text classification data set, the repository does contain [^94] https://github.com/Buddhilive/sltk/blob/main/data/sin.csv a data set of 26k Sinhala news titles with classification labels. A visual Q&A benchmark data set named ALM-Bench [^95] https://huggingface.co/datasets/MBZUAI/ALM-Bench was created by Vayani et al. [164] to evaluate Q&A performance of LLMs on 100 languages including Sinhala. They initially machine translated LLaVA-Bench (In-the-Wild) dataset [165] data set and then extended it by searching for images under country name, language name, and cultural category on the web so that they can include culturally relevant questions and answer for each language. In their own experiment results, they report that GPT-4o [166] performance on Sinhala is superior to other tested LLMs on this data set. Ranathunga et al. [167] has created a text simplification dataset named SiTSE [^96] https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation. It has 1000 complex Sinhala texts along with two separate simplifications for each of them. Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data. Ranathunga et al. [168] created multiNER [^97] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil. It contains 99k tagged Sinhala words. They have used PER, ORG, LOC, and MISC tags as well as O. Singh et al. [169] introduced Global-MMLU [^98] https://huggingface.co/datasets/CohereForAI/Global-MMLU, a multilingual extension of the original English Massive Multitask Language Understanding (MMLU) dataset [170]. They claim to have improved the data set beyond what Üstün et al. [171] did before. Chavinda and Thayasivam [172] introduces two publicly available Sinhala hate speech data sets. One is from Facebook [^99] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter100100100https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech. A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released101101101https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [173]. One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish.
@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [174]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [175, 176, 177, 178, 179, 180] as well. The dictionary by Kulatunga [181] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [182] which bears a more than passing resemblance to the above dictionary by Kulatunga [181]Hettige and Karunananda [79] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [183]. The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github102102102https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara103103103http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [186] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [37] to information extraction based on semantic similarity [187, 188] or semantic oppositeness [189]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [48]. For a time it was hosted on [190] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [191] have cloned to use in their application uploaded on github104104104https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [192] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [48].
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [205], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [207] which then extended to a Hidden Markov Model (HMM) based approach [208] and an analysis of unknown words [209, 210]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [211]. A RESTFul PoS tagging web service created by Jayaweera and Dias [212] using the above research can still be accessed105105105http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [213]. The study by Kothalawala et al. [214] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [215] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [216] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [115]. The study by Sathsarani et al. [217] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.

### v25

@@ Abstract
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field. This manuscript is at version $5.7.0$.
@@ 1 Introduction
To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10, 11]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. An extremely simplified version of the Indo-European language tree focusing mainly on Sinhala and its close relatives is shown in Fig 1. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [12, 13, 14, 15, 16, 17, 18]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [19, 20]. Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [21]. An extremely simplified evolution tree of the Sinhala script is shown in Fig 2. In it, the Sinhala script is shown as a singular derivation from Brahmi directly appearing with the contemporary script. But this is a simplification. We show the actual evolutionary eras of the Sinhala script in Appendix A where the gradual change from Brahmi to the contemporary script is clearer. It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22]. An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [23] using the SinMin Sinhala corpus created by Upeksha et al. [24, 25] is given in Appendix B.
Figure 1: A portion of the Indo-European language tree with a focus on Sinhala and its close relatives. Only the Elu sub-tree is shown exhaustively. All other sub-trees only list some languages for representation purposes chosen randomly or otherwise. Some of the languages shown as Extinct have archaeological proof (eg: Prakrit) while others were retroactively reconstructed by linguists by observing the commonalities of languages that they begot (eg: Proto-Indo-Aryan). The status of some languages has spread beyond linguistics to be a social and political question. In this depiction, we have attempted to side with academic consensus when available and the majority when not. As an example of one contentions point on this diagram note the positioning of Pali. While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
Figure 2: A portion of the writing systems (scripts) tree with a focus on Sinhala and its close relatives. It is important to note that this is distinct from the language tree shown in Fig 1. The writing systems used by a language being related, not being related, or the degree to which they are related is not directly correlated to the relevant languages being related, not being related, or the degree to which they are related. As an example, note that the writing systems of Sinhala and Tamil are reasonably closely related. But Tamil is not shown on Fig 1 given that it does not belong to the Indo-European language family. It belongs to the Dravidian language family. A language that is shown as a single entity in Fig 1 may have multiple entries in this diagram. An example is the language shown as Hindustani splits into Hindi and Urdu where the former uses the Devanagari script and the latter a derivation of Arabic script. Conversely, languages that may be shown as distinct entities in the full language tree, may map to a singular entity in the script tree. An example is the fact that English which would go under Proto-Germanic and Italian which would go under Proto-Italic in and extended version of Fig 1, would both map to Latin in this diagram given that both of them use the Latin script.
@@
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [28] and Wijeratne et al. [35]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [45, 46] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [47], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.7.0$. The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR. Given that this work was initially published in 2019, when citing this work, it is recommended to cite the most recent version to avoid the reviewers questioning its continued contemporary relevance. An example of an updated reference is given in Code Block LABEL:lst:ref.
@@
journal={arXiv preprint arXiv:1906.02358v25},
@@
Figure 10 in Appendix C shows the probability of studies from the institutions to which the most prolific authors from Figure 9 are affiliated citing institutions of the same set. We observe both interesting and disturbing trends which we discuss in the figure caption. Figure 11 in Appendix C shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests, denoted by the subsection titles of the Section 3 of this paper. If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication. For example, the paper Building a wordnet for Sinhala [48] is counted for both Nisansa de Silva and Gihan Dias. If a single paper contributes to more than one research area, that paper is counted for all of the research areas to which it contributes. For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [28] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
The remainder of this survey is organized as follows: Section 2 introduces some important properties and conventions of the Sinhala language, which are important for the development and understanding of Sinhala NLP. Section 3 discusses the various tools and research available for Sinhala NLP. In this section, we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work. We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka. Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work. Finally, Section 5, concludes the survey.
@@ 3.2 Data Sets
Singh et al. [159] introduced the Aya Data set [^87] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM). They also introduced Aya Collection [^88] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages. Both of these data sets include Sinhala. Further, they also released Aya Evaluation suite [^89] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality. Kudugunta et al. [160] introduced MADLAD-400909090https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents. They have also released the trained checkpoints on GitHub [^91] https://github.com/google-research/google-research/tree/master/madlad_400. A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al. [161] along with data [^92] https://huggingface.co/datasets/afaji/cvqa and a leader board [^93] https://eval.ai/web/challenges/challenge-page/2305/overview. Pushpananda et al. [162] claims to have created a manually annotated parallel corpus named TamSiPara with 25k parallel sentences in the Si-Ta direction and 22k parallel sentences in the Ta-Si direction. But it is not freely and openly accessible. While the stated purpose of Ranasinghe [163] is creating a tokenizer for Sinhala and not creating a text classification data set, the repository does contain [^94] https://github.com/Buddhilive/sltk/blob/main/data/sin.csv a data set of 26k Sinhala news titles with classification labels. A visual Q&A benchmark data set named ALM-Bench [^95] https://huggingface.co/datasets/MBZUAI/ALM-Bench was created by Vayani et al. [164] to evaluate Q&A performance of LLMs on 100 languages including Sinhala. They initially machine translated LLaVA-Bench (In-the-Wild) dataset [165] data set and then extended it by searching for images under country name, language name, and cultural category on the web so that they can include culturally relevant questions and answer for each language. In their own experiment results, they report that GPT-4o [166] performance on Sinhala is superior to other tested LLMs on this data set. Ranathunga et al. [167] has created a text simplification dataset named SiTSE [^96] https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation. It has 1000 complex Sinhala texts along with two separate simplifications for each of them. Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data. Ranathunga et al. [168] created multiNER [^97] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil. It contains 99k tagged Sinhala words. They have used PER, ORG, LOC, and MISC tags as well as O. Singh et al. [169] introduced Global-MMLU [^98] https://huggingface.co/datasets/CohereForAI/Global-MMLU, a multilingual extension of the original English Massive Multitask Language Understanding (MMLU) dataset [170]. They claim to have improved the data set beyond what Üstün et al. [171] did before. Chavinda and Thayasivam [172] introduces two publicly available Sinhala hate speech data sets. One is from Facebook [^99] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter100100100https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech. A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released101101101https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [173]. One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish. Burchell et al. [174] introduced HPLT v2102102102https://huggingface.co/datasets/HPLT/HPLT2.0_cleaned a multilingual monolingual and parallel corpora that includes 273k Sinhala Sentences. These sentences were extracted from the Internet Archive(IA) and the Common Crawl (CC). They have also publicly released their translation models103103103https://huggingface.co/collections/HPLT/hplt-20-uni-direction-translation-models-67f2fc7ae54845f9b182957a.
@@ 3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [175]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [176, 177, 178, 179, 180, 181] as well. The dictionary by Kulatunga [182] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [183] which bears a more than passing resemblance to the above dictionary by Kulatunga [182]Hettige and Karunananda [79] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [184]. The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github104104104https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara105105105http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
@@ 3.4 WordNets
WordNets [187] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [37] to information extraction based on semantic similarity [188, 189] or semantic oppositeness [190]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [48]. For a time it was hosted on [191] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [192] have cloned to use in their application uploaded on github106106106https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [193] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [48].
@@ 3.6 Part of Speech Taggers
Within a single group which did one of the above stochastic studies [206], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [208] which then extended to a Hidden Markov Model (HMM) based approach [209] and an analysis of unknown words [210, 211]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [212]. A RESTFul PoS tagging web service created by Jayaweera and Dias [213] using the above research can still be accessed107107107http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [214]. The study by Kothalawala et al. [215] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [216] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [217] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [115]. The study by Sathsarani et al. [218] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.